what is the magnetic field at the center of n sided regular polygon (derivation)?

Grade:12

2 Answers

Nitesh
6 years ago
A polygon which have n shaped is likely to be a circle Assume a n side polygon is hexagon whose two diagonal makes 2π/4 so n side polygon has angle 2π/n.Make a straight line from centre to side perpendicularly which bisects 2π/n in 2parts of π/n.And half of its diagonal is R .so perpendiculare bisector length is Rcosπ/n.By using formula of magnetic field by straight wire the m.f(magnetic field) by only one side is dB=(u0* I (sinπ/n+sinπ/n))/2πRcosπ/n dB=(u0*I *sinπ/n)/πRcosπ/nM.F by n sides dB*n="(u0*I * sinπ/n)n/πRcosπ/n."
ARYAN MATHUR
5 years ago
It is simply NB because the theta at each case is 90° ,also with the number of sides there is a equal distance from center, so the answer i get is NB.